OpenStudy (anonymous):

I need help!!! please A piece of rope is long enough to go around a circular flower-bed 5 times. What is the radius and area of the flower-bed if the rope is 660m in length?

OpenStudy (rulnick):

circumf = 660/5 m = 132 m radius * 2 * pi = circumf = 132 m radius = 66/pi m area = pi (radius)^2 = pi (66/pi m)^2 = 66^2/pi m

OpenStudy (anonymous):

the circumfrance of a circle is (2)(pi)(r). you want to go around your circle 5 times so will need to multiply that formula by 5. So you'll have (10)(pi)(r). Set that equal to your length of 660. And solve.

OpenStudy (anonymous):

ok it looks very confusing and messy rulnick

OpenStudy (rulnick):

radius = 21.01 m area = 1387 m^2

OpenStudy (anonymous):

\[10(\pi)(r)=660\]\[r=660/10\pi\] \[r=21.199\]